OpenStudy (anonymous):

LMNO is a parallelogram. If NM= x+ 19 and OL=3x+7. Find the value of x, then find NM and OL

OpenStudy (callisto):

NM = OL x+19 = 3x+7 2x = 12 x = 12/2 =?

OpenStudy (callisto):



OpenStudy (anonymous):

a) x=8, NM=27, OL=27 b) x=8, NM=25, OL=27 c) x=6, NM=25, OL=25 d) x=6, NM= 27, OL=25

OpenStudy (callisto):

NM= x+ 19 = ? +19 = .. OL=3x+7 = 3 (?) +7 =... Where ? = 12/2

OpenStudy (anonymous):

Where did you get 12 from?

OpenStudy (callisto):

Please see my first comment NM = OL (prop. of //gram) x+19 = 3x+7 2x = 12 x = 12/2 =?

OpenStudy (anonymous):

Oh nvm.. x+19=3x -x 19=2x+7 -7 -7 ___________ 12 = 2x x=6

OpenStudy (anonymous):

&& the sides would equal, so it would be 25 & 25, not 25 & 27.

OpenStudy (anonymous):

Correct?

OpenStudy (callisto):

Yes :)